OpenStudy (anonymous):

y = x -1 tangent to the graph of f(x) at f'(-3)

OpenStudy (anonymous):

I need to find the deriative: Is f(x) = -3 y = x - 1 so x-(-3) - 1 = -4 so y = 4 x = -3 My point is (-3, -4) Do I use the point slope equation y - (-4) = -3 (x - 3) I get y = -3x + 5 How do I find the derivative of f(x) = -3